OpenStudy (anonymous):

suppose y varies directly with x and y=12 when x=-3 what is the value of y when x=6

OpenStudy (alexandervonhumboldt2):

y/x=y/x y/6=12/(-3) y/6=-4 y=-24

OpenStudy (alexandervonhumboldt2):

use ratios

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wwhat

OpenStudy (tkhunny):

"y varies directly with x" MEANS y = kx Fill in the given values and solve for k.